# Are AI Detectors Good Enough? A Survey on Quality of Datasets With Machine-Generated Texts

### v2

@@ Abstract
Are AI Detectors Good Enough? A Survey on Quality of Datasets With Machine-Generated Texts
German Gritsaithanks: Equal contributions Advacheck OÜ, Tallinn, Estonia Université Grenoble Alpes, Grenoble, France gritsai@advacheck.com                        Anastasia Voznyuk Advacheck OÜ, Tallinn, Estonia voznyuk@advacheck.com
Andrey Grabovoy Advacheck OÜ, Tallinn, Estonia grabovoy@advacheck.com                        Yury Chekhovich Advacheck OÜ, Tallinn, Estonia chekhovich@advacheck.com
@@ 1 Introduction
The quality of large language models (LLMs) has grown tremendously in the last five years, making their output almost indistinguishable from human-written texts  Chang et al. (2024). This expanded the application fields of these models, as many routine tasks can be entrusted to them nowadays. However, they can be used for creating texts that are intended to be written and fact-checked by humans. Example of such misuse is generation of fake news Zellers et al. (2019); Zhou et al. (2023), which can mislead readers of generated content. Teachers raise another concern, as many students complete assignments with LLMs Koike et al. (2024); Ma et al. (2023), undervaluing the purpose of the educational process. Machine-generated fragments also appear in academic articles more often with the growth of chatbots and reach several tens of percent  Liang et al. (2024); Gritsay et al. (2023a). More than 60,000 scientific papers in the last year alone contained the evidence of the use of machine generation Gray (2024). All of that prove that it is crucial to develop systems able to counter the misuse of artificial data and signal to the reader that the content they read is generated.
@@
Another concern is the Web, overflowing with machine-generated content, often of poor quality. Such texts contribute bias to publicly available texts on the Internet, through false facts, hallucinations and spelling errors. Given the current agenda of using texts from the Internet to train new language models, all this bias will be inadvertently added to the model. Moreover,  Villalobos et al. (2022) revealed that the human-written data will run out by 2028. That means that the training sets for language models in the future will include a large amount of generated content. Such self-consuming will result in the substantial degradation of the model’s abilities Alemohammad et al. (2023). Furthermore, the trend is evolving in such a way that human-written texts on almost any topic will be much harder to retrieve. While for texts dated even 5 years ago we are confident as the usage of generation was extremely rare, we cannot state the same for more recent texts.
@@
Therefore, detectors, capable to distinguish human-written texts from AI-generated texts, and whose detection quality can be guaranteed, are necessary for many fields. We believe one of the key factors for building reliable detectors is the high-quality artificial text collections that can be used for training and evaluation. In this paper, we would like to estimate the quality of the available generated texts from competitions and research papers. Sometimes we see that some methods from participants of the competitions reach almost perfect (up to 99.9%) metric score, meanwhile in the wild we observe a noticeable decline in performance. Such results look confusing, because the models become more and more advanced, seemingly making the detecting task more challenging, meanwhile participants of competitions still reach almost perfect scores, bringing up the question about quality of generated data in the provided datasets. Are the devised methods really good or is the data easy enough for detectors to solve the seemingly hard detection task?
Our contributions are as follows.
@@ 2.1 The Task of AI-generated text detection
The approach to tackle the classification problem is to utilize linguistic, stylometric, and statistical features for classifiers Jawahar et al. (2020); Fröhling and Zubiaga (2021). Some internal metrics, such as perplexity or its modifications Hans et al. (2024) can also be helpful as a inference method. Also, perturbing texts can also provide valuable information, as described in DetectGPT Mitchell et al. (2023) method, where log-probabilities between between original and modified texts are compared. Finally, finetuning encoder-based model, like DeBERTa He et al. (2021), is considered the best solution for this task Uchendu et al. (2021); Macko et al. (2023).
@@ 2.2 Evaluating Generated Text
As for evaluating the quality of the generated data itself, it has become more common to evaluate it with the help of LLMs Xu et al. (2023). This approach does not require any human reference, unlike ROUGE Lin (2004). However, the downside is the output of model-evaluator needs to be unified, not always interpretable and model-evaluator scores can be skewed. Alternative approach is suggested by  Zhu and Bhat (2020), where the text is evaluated based on several linguistic criteria, such as grammar or coherence.
@@ 2.3 Datasets with Artificial Content
There are a number of surveys about machine-generated content detection with an overview of the datasets Jawahar et al. (2020); Wu et al. (2023), however, few works focus on the quality of data in the available datasets, despite it being an important aspect of the task. Building AI-generated content detectors requires high-quality labelled data that involve substantial financial, computational, and human resources. The human evaluators should check that the dataset does not contain corrupted generations, that the texts are coherent and grammatically correct. We will describe the datasets we used in our analysis and experiments in Section 3.
@@ 2.4 Shared Tasks on AI-Generated content Detection
Shared tasks help to advance research on detecting AI-generated content, offering new variations on the tasks and data for evaluation, encouraging participants to come up with novel ideas for detectors robust to the change of language, domain, or generating model. Participants explore approaches ranging from transfer learning on complex text features to utilising and fine-tuning LLMs for these tasks (2022; 2022; 2023b; 2024; 2024; 2024; 2024; 2024). These efforts have highlighted challenges such as handling multilingual data and adapting to rapidly evolving generative models. Some participants also provide some analysis of the given data or even discuss some flaws with generated texts in it Voznyuk and Konovalov (2024).
@@ 3 Datasets
Dataset Year Language # Texts, $10^{3}$ Texts, G / H, $10^{3}$ Average Length, G / H Median Length, G / H
GPT2 2019 en 1250 1000 / 250 2941 / 2616 3245 / 2459
TweepFake 2019 en 20.7 10.4 / 10.4 104 / 118 89 / 94
HC3 2023 en, zh 85.4 26.9 / 58.5 1011 / 681 1012 / 422
GhostBuster 2023 en 21 18 / 3 3345 / 3391 3440 / 2911.5
MGTBench 2024 en 23.7 20.7 / 3 1596 / 3391 1226 / 2911.5
MAGE 2024 en 436 152.3 / 284.2 1139 / 1282 706 / 666
M4 2024 en, zh, ru, bg, ur, id 89.5 44.7 / 44.7 1588 / 3162 1454 / 1697
OutFox 2024 en 57.6 43.2 / 14.4 2686 / 2238 2311 / 1992
DAGPap22 2022 en 5.3 3.6 / 1.6 799 / 1180 680 / 1126.5
RuATD 2022 ru 129 64.5 / 64.5 237 / 221 99 / 95
AuTex 2023 en, es 65.9 33.1 / 32.8 315 / 297 386 / 351
IberAuTex 2024 es, en, ca, gl, eu, pt 98 52.5 / 45.4 1037 / 1058 981 / 1018
PAN24 2024 en 15.2 14.1 / 1.1 2641 / 3007 2731 / 2868
SemEval24 Mono 2024 en 34.2 18 / 16.2 2465 / 2358 2570 / 2083.5
SemEval24 Multi 2024 en, ar, de, it 42.3 22.1 / 20.2 2218 / 2257 2270 / 2032
MGT-1 Mono 2025 en 610.7 381.8 / 228.9 1448 / 1541 1208/ 1080
MGT-1 Multi 2025 en, zh, it, ar, de, ru, bg, ur, id 674 416.1 / 257.9 1423 / 1445 1195 / 1032
@@ 3.1 Datasets From Shared Tasks
  1. 1.
    DAGPap 2022 Kashnitsky et al. (2022) introduced a dataset of human- and machine-written scientific excerpts collected by Elsevier.
  2. 2.
    RuATD 2022 Shamardina et al. (2022) focused on human- and machine-written documents in Russian, covering a wide range of themes.
  3. 3.
    AuTexTification 2023 Sarvazyan et al. (2023) provided texts in English and Spanish, covering five distinct domains.
  4. 4.
    IberAuTexTiification 2024 Sarvazyan et al. (2024) expanded on the previous competition with a multilingual (six Iberian languages), multi-domain, and multi-model focus.
  5. 5.
    Voight-Kampff Generative AI Authorship Verification 2024 Ayele et al. (2024), hereafter referred to as PAN 2024, tasked participants with identifying the human-authored text from two samples – one human-written and one machine-generated.
  6. 6.
    SemEval 2024 Task 8 Wang et al. (2024c) addressed domain, generator, and language shifts in generated texts. Training data included multiple languages (e.g., Chinese, Urdu, and Russian), but the test set was limited to English, Italian, German, and Arabic.
  7. 7.
    MGT Detection Task 1 (COLING 2025) Wang et al. (2025) was built on SemEval 2024 Task 8 by incorporating data generated by novel LLMs and expanding multilingual coverage of train and test sets.
3.2 Datasets from Papers
@@
HC3 (Human Chatbot Conversations Corpus) Su et al. (2024) features conversations between humans and chatbots, primarily used for research on chatbot responses and human-AI interaction analysis. This dataset is available for both English and Chinese, but we have focused only on the former.
GhostBuster Verma et al. (2024) aimed at detecting AI-generated content by comparing it to human-written text, often used in the context of identifying machine-generated misinformation or spam.
MGTBench (Machine Generated Text Benchmark) He et al. (2023) is a benchmark dataset designed to evaluate the quality of machine-generated text across various tasks, including fluency, coherence, and creativity.
MAGE (Model Augmented Generative Evaluation) Li et al. (2024) evaluates the performance of generative models by comparing outputs with human annotations, aiding in the development of more accurate generative AI models.
M4 (Multilingual, Multimodal, Multitask, Massive Dataset) Wang et al. (2024b) is a large-scale dataset designed for training models that can handle multiple languages, tasks, and modalities, making it useful for developing versatile AI systems. Although it is multilingual, we sampled only English texts.
TweepFake Fagni et al. (2021) contains real tweets written by humans and synthetic tweets, generated by various AI models, from bots, imitating human users.
Outfox Koike et al. (2024) contains triplets of essay problem statements, human-written essays, and LLM-generated essays. The students who wrote the essays range from 6th to 12th grade in the USA.
@@ 4.1 Baselines
We used several common classifiers as baselines on each dataset to obtain first-hand understanding of the dataset. The first one is mDeBERTa He et al. (2021), which is the current state-of-the-art model for multilingual machine-generated text detection Macko et al. (2023). The next baseline is DetectGPT framework Mitchell et al. (2023). We used GPT-2 Radford et al. (2019) as the base model and T5-Large Raffel et al. (2019) as perturbations generator. However, as DetectGPT requires intensive computational costs, we utilized Fast-DetectGPT Bao et al. (2024) that substitutes DetectGPT’s perturbation step with a more efficient sampling step. Finally, we used Binoculars Hans et al. (2024) as a method with improved perplexity score. The last two baselines require no fine-tuning which is crucial for detection task, as it is infeasible to train the detector on all existing artificial texts.
@@ 4.2 Topological Statistics
It was shown in Tulchinskii et al. (2023) that if we take the inner dimensionality of the manifold on the set of embeddings, we could separate human-written texts from machine-generated ones. The authors used persistence homology dimension (PHD) and showed that statistically human-generated texts have higher PHD than machine-generated texts, therefore introducing a novel detector. We calculated PHD on each set of texts. Additionally, in  Kushnareva et al. (2024) it was suggested to calculate PHD within sliding window. These intrinsic dimensions of the text within sliding window can be used as a feature for detector. The authors conclude that this metric as well can help to differentiate texts of different origin. To be able to compare datasets between each other, we came up with a score, utilizing KL-divergence, but we also needed it to be symmetrical. Let $h_{d}$, $m_{d}$ be distributions of intrinsic dimensions for two types of texts from the same dataset, of human and machine origin, then our $\text{KL}_{\text{TTS}}$ is following:
@@ 4.3 Perturbations and Shuffling
Based on the results of the text modification studies Sadasivan et al. (2024); Mitchell et al. (2023), which show how small perturbations affect machine reading comprehension systems, we decided to consider this way of possibly assessing the quality of a dataset. The key idea here is that AI models are sensitive to such adversarial changes, unlike humans. We considered two modification ideas: Adversarial Token Perturbation and Sentence Shuffling.
Adversarial Token Perturbation. In this approach we divide the text into tokens and randomly replace the token with a synonym from the WordNet collection Miller (1994) with a probability of 70%. We apply such a technique to each represented class. Using an encoder model, we obtain embeddings for each of the texts in the current dataset. Finally, we measure the average embedding shifts for the classes of human and generated texts. We obtain the embedding shifts using the cosine distance between the embeddings of the original texts and the modified ones. As a result, after modifications we obtain $\Delta_{\text{shift}}$ — the log difference of the average embedding shifts.
@@
where $n$ and $m$ – number of samples in the human and generated parts of the dataset respectively, $h_{h_{i}}^{o}$ – embedding of $i$-th fragment of human part of data, $h_{h_{i}}^{p}$ – the same embedding after perturbation. The same goes for $h_{m_{i}}^{o}$ and $h_{m_{i}}^{o}$ for machine-generated part of data. Finally, $\text{cos}_{d}$ is a function that measures the cosine distance between two vectors.
Sentence Shuffling. In this approach, we randomly swap sentences, thereby affecting the cohesion of the text. We try to find out the effect of artificial origin on the difference between the distributions after permutations. By dividing a fragment into sentences and randomly reversing the order of 70% of the selected sentences, we apply this technique to each represented class. Further, using the text encoding model, we obtain embeddings for each of the texts of the current dataset. Finally, we measure embedding shifts for the class of human and generated texts, and after that we convert the shifts into probability-like distributions. This allows us to obtain at the end $\text{KL}_{\text{shuffle}}(H,M)$ — the KL-divergence between the shifts of human and generated texts.
@@
$H(i)=\frac{\text{cos}_{d}(h_{h_{i}}^{o},h_{h_{i}}^{p})+\epsilon}{\sum_{j}\left(\text{cos}_{d}(h_{h_{j}}^{o},h_{h_{j}}^{p})+\epsilon\right)},$
and $M(i)$ has the same structure, except that instead of human class texts the generated class texts are used, $\epsilon$ is a small constant added to avoid division by zero.

### v3

@@ 1 Introduction
The quality of large language models (LLMs) has grown tremendously in the last five years, making their output almost indistinguishable from human-written texts  (Chang et al. 2024). This expanded the application fields of these models, as many routine tasks can be entrusted to them nowadays. However, they can be used for creating texts that are intended to be written and fact-checked by humans. An example of such misuse is the generation of fake news (Zellers et al. 2019; Zhou et al. 2023), which can mislead readers of such generated content. Teachers raise another concern, as many students complete assignments with LLMs (Koike, Kaneko, and Okazaki 2024; Ma et al. 2023), undervaluing the purpose of the educational process. Machine-generated fragments also appear in academic articles more often with the growth of chatbots and reach several tens of percent  (Liang et al. 2024; Gritsay et al. 2023a). More than 60,000 scientific papers in the last year alone contained evidence of the use of machine generation (Gray 2024). All of that proves that it is crucial to develop systems able to counter the misuse of artificial data and signal to the reader that the content they read is generated.
@@
Another concern is the Web, overflowing with machine-generated content, often of poor quality. Such texts contribute bias to publicly available texts on the Internet, through false facts, hallucinations and spelling errors. Given the current agenda of using texts from the Internet to train new language models, all this bias will be inadvertently added to the model. Moreover,  (Villalobos et al. 2022) revealed that the human-written data will run out by 2028. That means that the training sets for language models in the future will include a large amount of generated content. Such self-consuming will result in the substantial degradation of the model’s abilities (Alemohammad et al. 2023). Furthermore, the trend is evolving in such a way that human-written texts on almost any topic will be much harder to retrieve. While for texts dated even 5 years ago we are confident as the usage of generation was extremely rare, we cannot state the same for more recent texts.
@@
Therefore, detectors capable of distinguishing human-written texts from AI-generated texts and whose detection quality can be guaranteed, are necessary for many fields. We believe that one of the key factors for building reliable detectors is the high-quality artificial text collections that can be used for training and evaluation. In this paper, we would like to estimate the quality of the available generated texts from competitions and research papers. Sometimes we see that some methods from participants of the competitions reach almost perfect (up to 99.9%) metric score, meanwhile in the wild we observe a noticeable decline in performance. Such results look confusing, because the models become more and more advanced, seemingly making the detecting task more challenging, meanwhile participants of competitions still reach almost perfect scores, bringing up the question about quality of generated data in the provided datasets. Are the devised methods really good or is the data easy enough for detectors to solve the seemingly hard detection task?
Our contributions are as following
@@ 2.1 The Task of AI-generated text detection
The first approaches to tackle the classification problem were to utilise some linguistic, stylometric, and statistical features for classifiers (Jawahar, Abdul-Mageed, and Lakshmanan 2020; Fröhling and Zubiaga 2021). However, while these methods performed well for the texts from the first language models, nowadays models are advanced enough to output texts that are almost indistinguishable from human-written ones, therefore these methods are currently not reliable enough. The next category are zero-shot methods that employ metrics, such as perplexity or its modifications (Hans et al. 2024), which can be helpful as an inference method when training is not available. Another approach that do not require training is perturbing texts, which can also provide valuable information. For example, one can compare log-probabilities between between original and perturbed texts, as described in DetectGPT (Mitchell et al. 2023) method. Finally, methods based on fine-tuning encoder-based models, such as DeBERTa (He et al. 2021), are currently considered the state-of-the-art approach for the detection task (Uchendu et al. 2021; Macko et al. 2023).
@@ 2.2 Evaluating Generated Text
As for evaluating the quality of the generated data itself, it has become more common to evaluate it with the help of LLMs (Xu et al. 2023). This approach does not require any human reference, unlike ROUGE (Lin 2004). However, the output of model-evaluator needs to be unified, is not always interpretable, and model-evaluator scores can be skewed. Alternative approach is suggested by  (Zhu and Bhat 2020), where the text is evaluated based on several linguistic criteria, such as grammar or coherence.
@@ 2.3 Datasets with Artificial Content
There are a number of surveys of machine-generated content detection with an overview of the datasets (Jawahar, Abdul-Mageed, and Lakshmanan 2020; Wu et al. 2023), however, few works focus on the quality of data in the available datasets, despite it being an important aspect of the task. Building AI-generated content detectors requires high-quality labelled data that involve substantial financial, computational, and human resources. The human evaluators should check that the dataset does not contain corrupted generations, that the texts are coherent and grammatically correct. We will describe the datasets we used in our analysis and experiments in Section 3.
@@ 2.4 Shared Tasks on AI-Generated content Detection
Shared tasks advance research on detecting AI-generated content forward by offering new variations on tasks and providing data for evaluation, encouraging participants to come up with novel ideas for detectors robust to the change of language, domain, or generating model. Participants explore approaches ranging from transfer learning on complex text features to utilising and fine-tuning LLMs for these tasks (2022; 2022; 2023b; 2024; 2024; 2024; 2024; 2024). These efforts have highlighted challenges such as handling multilingual data and adapting to rapidly evolving generative models. Some participants also provide some analysis of the given data or even discuss some flaws with the generated texts (Voznyuk and Konovalov 2024).
@@ 3 Datasets
Dataset Language Num. of Texts, $10^{3}$ Num. of Texts, G / H, $10^{3}$ Average Length, G / H Median Length, G / H
GPT2 en 1250 1000 / 250 2941 / 2616 3245 / 2459
TweepFake en 20.7 10.4 / 10.4 104 / 118 89 / 94
HC3 en, zh 85.4 26.9 / 58.5 1011 / 681 1012 / 422
GhostBuster en 21 18 / 3 3345 / 3391 3440 / 2911.5
MGTBench en 23.7 20.7 / 3 1596 / 3391 1226 / 2911.5
MAGE en 436 152.3 / 284.2 1139 / 1282 706 / 666
M4 en, zh, ru, bg, ur, id 89.5 44.7 / 44.7 1588 / 3162 1454 / 1697
OutFox en 57.6 43.2 / 14.4 2686 / 2238 2311 / 1992
DAGPap22 en 5.3 3.6 / 1.6 799 / 1180 680 / 1126.5
RuATD ru 129 64.5 / 64.5 237 / 221 99 / 95
AuTex en, es 65.9 33.1 / 32.8 315 / 297 386 / 351
IberAuTex es, en, ca, gl, eu, pt 98 52.5 / 45.4 1037 / 1058 981 / 1018
PAN24 en 15.2 14.1 / 1.1 2641 / 3007 2731 / 2868
SemEval24 Mono en 34.2 18 / 16.2 2465 / 2358 2570 / 2083.5
SemEval24 Multi en, ar, de, it 42.3 22.1 / 20.2 2218 / 2257 2270 / 2032
MGT-1 Mono en 610.7 381.8 / 228.9 1448 / 1541 1208/ 1080
MGT-1 Multi en, zh, it, ar, de, ru, bg, ur, id 674 416.1 / 257.9 1423 / 1445 1195 / 1032
@@ 3.1 Datasets From Shared Tasks
  • DAGPap 2022 (Kashnitsky et al. 2022) introduced a dataset of human- and machine-written scientific excerpts collected by Elsevier.
  • RuATD 2022 (Shamardina et al. 2022) focused on human- and machine-written documents in Russian, covering a wide range of themes.
  • AuTexTification 2023 (Sarvazyan et al. 2023) provided texts in English and Spanish, covering five distinct domains.
  • IberAuTexTiification 2024 (Sarvazyan et al. 2024) expanded on the previous competition with a multilingual (six Iberian languages), multi-domain, and multi-model focus.
  • Voight-Kampff Generative AI Authorship Verification 2024 (Ayele et al. 2024), hereafter referred to as PAN 2024, tasked participants with identifying the human-authored text from two samples – one human-written and one machine-generated.
  • SemEval 2024 Task 8 (Wang et al. 2024c) addressed domain, generator, and language shifts in generated texts. Training data included multiple languages such as Chinese, Urdu, and Russian, but the test set was limited to English, Italian, German, and Arabic.
  • MGT Detection Task 1 (COLING 2025) (Wang et al. 2025) was built on SemEval 2024 Task 8 by incorporating data generated by novel LLMs and expanding the multilingual coverage of the train and test sets.
3.2 Datasets from Research Papers
@@
  • HC3 (Human Chatbot Conversations Corpus) (Su et al. 2024) features conversations between humans and chatbots, primarily used for research on chatbot responses and human-AI interaction analysis. This dataset is available for both English and Chinese, but we have focused only on the former.
  • GhostBuster (Verma et al. 2024) aimed at detecting AI-generated content by comparing it to human-written text, often used in the context of identifying machine-generated misinformation or spam.
  • MGTBench (Machine Generated Text Benchmark) (He et al. 2023) is a benchmark dataset designed to evaluate the quality of machine-generated text across various tasks, including fluency, coherence, and creativity.
  • MAGE (Model Augmented Generative Evaluation) (Li et al. 2024) evaluates the performance of generative models by comparing outputs with human annotations, aiding in the development of more accurate generative AI models.
  • M4 (Multilingual, Multimodal, Multitask, Massive Dataset) (Wang et al. 2024b) is a large-scale dataset designed for training models that can handle multiple languages, tasks, and modalities, making it useful for developing versatile AI systems. Although it is multilingual, we sampled only English texts.
  • TweepFake (Fagni et al. 2021) contains real tweets written by humans and synthetic tweets, generated by various AI models, from bots, imitating human users.
  • Outfox (Koike, Kaneko, and Okazaki 2024) contains triplets of essay problem statements, human-written essays, and LLM-generated essays. The students who wrote the essays range from 6th to 12th grade in the USA.
@@ 4.1 Baselines
In Section 2.1 we described three main categories of methods for tackling the detection task. We chose a method from each category, that served as a baseline to obtain first-hand understanding of each dataset. For the perturbation-based methods we used DetectGPT framework with GPT-2 (Radford et al. 2019) as the base model and T5-Large (Raffel et al. 2019) as perturbations generator. However, due to intensive computational costs of DetectGPT, we utilised Fast-DetectGPT (Bao et al. 2024) that substitutes DetectGPT’s perturbation step with a more efficient sampling step. For the zero-shot methods we used Binoculars (Hans et al. 2024) with improved perplexity score. These two baselines need no fine-tuning, which is an important aspect for detection task, as it is infeasible to train the detector for every domain and generator. Lastly, as encoder-based method we used mDeBERTa (He et al. 2021), which is the current state-of-the-art model for multilingual machine-generated text detection (Macko et al. 2023). By taking these three detectors, we covered all main categories of detectors.
@@ 4.2 Topological Statistics
It was shown in (Tulchinskii et al. 2023) that if we take the inner dimensionality of the manifold on the set of embeddings, we could separate human-written texts from machine-generated ones. The authors used persistence homology dimension (PHD) and showed that statistically human-generated texts have higher PHD than machine-generated texts, therefore introducing a novel detector. We calculated PHD on each set of texts. Additionally, in  (Kushnareva et al. 2024) it was suggested to calculate PHD within sliding window. These intrinsic dimensions of the text within sliding window can be used as a feature for detectors. The authors show that the metric is robust to the change of domain and generators. To be able to compare datasets between each other, we came up with a symmetrical score, utilising KL-divergence. Let $h_{d}$, $m_{d}$ be distributions of intrinsic dimensions for two types of texts from the same dataset, of human and machine origin, then our $\text{KL}_{\text{TTS}}$ is following:
@@ 4.3 Perturbations and Shuffling
Based on the results of the text modification studies (Sadasivan et al. 2024; Mitchell et al. 2023), which show how small perturbations affect machine reading comprehension systems, we decided to consider this way of possibly assessing the quality of a dataset. The key idea here is that AI models are sensitive to such adversarial changes, unlike humans. We considered two modification ideas: Adversarial Token Perturbation and Sentence Shuffling.
Adversarial Token Perturbation. In this approach we divide the text into tokens and randomly replace the token with a synonym from the WordNet collection (Miller 1994) with a probability of 70%. We apply such a technique to each represented class. Using an encoder model, we obtain embeddings for each of the texts in the current dataset. Finally, we measure the average embedding shifts for the classes of human and generated texts. We obtain the embedding shifts using the cosine distance between the embeddings of the original texts and the modified ones. As a result, after modifications we obtain $\Delta_{\text{shift}}$ — the log difference of the average embedding shifts.
@@
where $n$ and $m$ are number of samples in the human and generated parts of the dataset respectively, $h_{h_{i}}^{o}$ – embedding of the $i$-th fragment of human part of data, $h_{h_{i}}^{p}$ – the same embedding after perturbation. Similarly, $h_{m_{i}}^{o}$ and $h_{m_{i}}^{o}$ are embeddings for machine-generated texts. Finally, $\text{cos}_{d}$ is a function that measures the cosine distance between two vectors.
Sentence Shuffling. In this approach, we randomly swap sentences, thereby affecting the cohesion of the text. We try to find out the effect of artificial origin on the difference between the distributions after permutations. By dividing a fragment into sentences and randomly reversing the order of 70% of the selected sentences, we apply this technique to each represented class. Then, using the text encoding model, we obtain embeddings for each of the texts from the current dataset. Finally, we measure embedding shifts for the class of human and generated texts, and after that we convert the shifts into probability-like distributions. This allows us to obtain at the end $\text{KL}_{\text{shuffle}}(H,M)$ — the KL-divergence between the shifts of human and generated texts.
@@
$H(i)=\frac{\text{cos}_{d}(h_{h_{i}}^{o},h_{h_{i}}^{p})+\epsilon}{\sum_{j}\left(\text{cos}_{d}(h_{h_{j}}^{o},h_{h_{j}}^{p})+\epsilon\right)}.$
$M(i)$ has the same structure as $H(i)$, except that instead of human class texts the generated class texts are used, $\epsilon$ is a small constant added to avoid division by zero.